# Laser interferometry as a robust neuromorphic platform for machine learning

### v2

@@ IV.2.1 Binary Classification
We generated data points for the last two problems using scikit-learn with the same Gaussian noise as in the XOR case. We generated 100 data points and used a 50/50 split of the data into training and test sets for all problems. The hyperparameters are given in Table 2. We used a three-mode circuit with three layers for all binary classification problems and encode the inputs into the first two modes with the remaining mode being vacuum. The output logits were taken to be the expectation of the $Q$ quadrature of the first two modes of the circuit.
Hyperparameter Value
Learning rate ($\gamma$) $10^{-2}$
Learning rate decay ($\chi$) $0.999$
11st moment estimates decay rate ($\beta_{1}$) $0.9$
22nd moment estimates decay rate ($\beta_{2}$) $0.999$
Number of epochs $1000$
Batch size $32$
Table 2: Hyperparameters used when training the optical circuit for binary classification tasks.
The training data for each of the problems are shown in Fig.4.
@@
Fig.6 shows the performance of the model when trained with no photon loss and with photon loss where $\eta=0.5$ for the concentric circles problem. We can see that, as in the regression case, the model could still learn an appropriate nonlinear decision boundary in the presence of high losses. From the accompanying heatmaps we observed that, similar to the regression case larger displacements were applied in the lossy case. However, in contrast to the regression example all displacements throughout the network were higher in the lossy case, not just for the modes and layers closer to the output homodyne measurements.
@@
IV.2.2 Multilabel Classification
In addition to the previous sections which demonstrate this ONN model can indeed solve non-trivial nonlinear learning problems we also investigate the performance of our network on more challenging and commonly used benchmarks of multilabel classification. We used three examples of multilabel classification: the Iris data set which is a commonly used benchmark [3, 2, 59, 30, 23, 57], the handwritten digits dataset considered in Refs. [50, 52] and the vowel recognition dataset which is widely used in benchmarking of physical neural networks [47, 42, 21, 55, 33, Bandyopadhyay2024, 36]. The first two datasets are loaded via scikit-learn.
@@ IV.2.2 Multilabel Classification
The Iris dataset consists of 150 samples from three different species (classes) of Iris flower (setosa, virginica, and versicolor) with each sample being represented as a four dimensional vector with elements corresponding to petal width, petal length, sepal width and sepal length in cm [16]. We used 75 random samples for training and the rest for testing and normalized each input feature to be on the range $[0,1]$. We used a four mode circuit with three layers, leading to 42 trainable parameters, and the same hyperparameters used in the binary classification section. We encode the four features into the four modes as done in our previous experiments. The expectation value of the $Q$ quadrature of the first three modes are taken to be our models output logits. Fig.8 shows the confusion matrix of the ONN evaluated on the test set.
Figure 8: Confusion matrix for ONN classifier on test set of the Iris dataset.
We can see that the test accuracy is $94.67\%$. The two species the model struggles slightly to differentiate are virginica and versicolor. This is in accordance with previous findings using this dataset in the literature [23] as the two species have significant overlap in feature space particularly with their sepal widths and sepal lengths as shown in the pair plot in Fig.9.
Hyperparameter Value
Learning rate ($\gamma$) $10^{-2}$
Learning rate decay ($\chi$) $0.999$
11st moment estimates decay rate ($\beta_{1}$) $0.9$
22nd moment estimates decay rate ($\beta_{2}$) $0.999$
Number of epochs $500$
Batch size $512$
Table 3: Hyperparameters used when training the optical circuit for the digits classification task.
Figure 9: Pair plot of Iris dataset.
Figure 10: Examples of handwritten digits and their corresponding ground truth labels.
@@
The handwritten digits dataset consists of 1797 gray scaled images of handwritten digits from 0 to 9 and was considered. For visualization purposes we show five example images in Fig.10. We used 60% of the data (1078 images) for training and 40% (719 images) for testing. The pixel values were normalized to be between $[0,1]$ and the images were flattened into a $64$ dimensional vector. For this problem we used a 64 mode circuit with four layers, leading to $8576$ trainable parameters, and took the expectation value of the $Q$ quadrature of the first 10 modes as the output logits of the model. We encode the 64 normalized pixel intensities into the 64 modes of the optical circuit. The hyperparameters used to train the ONN for this task are shown in Table 3. We achieved a test accuracy of 98.19% as can be seen by examining the confusion matrix of the trained model in Fig.11.
@@
We consider the vowel recognition dataset from Refs. [42, 55, 33, 36], which is a subset of the dataset used in Ref. [19], that consists of pronunciations of seven vowels by 37 female speakers where each vowel is characterized by 12 different formant frequencies. Specifically, each data point in the dataset is represented as a 12-dimensional vector containing the first three formants $F_{1},F_{2}$ and $F_{3}$ sampled at the steady state, $20\%$, $50\%$, and $80\%$ of the vowels pronunciation duration respectively. Each speaker pronounces each of the seven vowels leading to a dataset with 259 data points. We use a 12 mode circuit with three layers, leading to 270 trainable parameters, and the same hyperparameters as in the binary classification section with the learning rate set to $10^{-3}$. We use a random split of the data into 129 training samples and 130 testing samples. We normalize the 12 features such that they are all between $[0,1]$. From the confusion matrix in Fig. 12 we can see that the ONN achieves a test accuracy of 98.46%.
@@
Figure 11: Confusion matrix for ONN classifier on test set of the handwritten digits dataset.
Figure 12: Confusion matrix for ONN classifier on test set of the vowel recognition dataset.
V Conclusion
We have proposed a method for constructing a linear optical neural network using laser interferometry and field displacements, which is the novelty claim of this paper. We demonstrated that this classical optical model displays remarkable resilience to photon loss which is a prominent source of imperfections in modern photonic integrated circuits [Bogaerts2020]. The linear optical character of the circuit also makes it highly suitable for in situ training using various existing training protocols based on parameter shift rules [45], physical backpropagation [20], physics-aware training [55] or forward-only training [33, 35, Bandyopadhyay2024, 57]. All this makes it particularly appealing for on-chip integrated optics implementations. We also investigated how our network performs on a variety of machine learning problems and its robustness to variance in the gradient estimates provided by parameter shift rules.
@@ V Conclusion
One straightforward quantum improvement would be to add an inline squeezer of the amplitude quadrature $Q$ before its measurement by homodyne detection during both training and inference. The corresponding quantum noise reduction will allow proportionally fewer measurements to determine the average $q=\langle Q\rangle$ consequently enabling faster inference and gradient estimation time. Note that just 10 dB of squeezing reduces the variance by ten times and that on-chip quantum homodyne detection has been demonstrated at a bandwidth of 9 GHz [Tasker2021] and PIN photodiode detection bandwidth has been demonstrated to reach 265 GHz [Lischke2021]. Another, albeit more complex, quantum improvement is swapping the single $M\times M$ interferometer with a set of $M$ single mode squeezers which are sandwiched between two $M\times M$ interferometers. This configuration, based on the singular value decomposition, is known to allow one to exactly mimic the functionality of fully connected linear layers within a classical neural network [25].
For future work one could explore how these networks can more closely approximate the expressivity of classical neural networks. For example, classical neural networks inherently have a feed forward structure in which outputs from one layer are composed with the nonlinear transformation (activation function) in the next layer leading to the network learning hierarchical and distributed representations of the inputs at each layer [8]. An interesting example of such feedforward (though a nonlinear optical one, using field encoding rather than parameter encoding) can be found in Ref. [Bandyopadhyay2024]. Furthermore, the existence of universal function approximation theorems for classical neural networks [14, 29] provides a strong theoretical basis for their expressivity and possible extensions of these theorems to our work merits further investigation. Lastly, extensions to implementing convolutional, recurrent and transformer layers are also important next steps.
@@
Acknowledgements.
We thank Benjamin Scellier, Maria Schuld and Catalina Albornoz for useful discussions. OP was supported by U.S. National Science foundation grants PHY-2514971 and ECCS-2530171. JMS acknowledges support from the National Science Foundation via DMR-2204312.
Appendix A Using intensity measurements
Many photonic neural networks take the intensity of the optical modes to be the network output and the efficient in situ training protocol proposed in Ref. [20] utilizes intensity measurements to extract gradients. Here we investigate the effect using intensity over $Q$ quadrature expectations has on our models performance. The intensity of an optical field is defined as $\langle N\rangle$, where $N=a^{\dagger}a$ is the photon number operator, and can be written in terms of the quadrature expectations as
$\langle N\rangle=\frac{1}{2}(\langle Q^{2}\rangle+\langle P^{2}\rangle-1)$ (24)
which for coherent states reduces to
$\langle N\rangle=\frac{1}{2}(\langle Q\rangle^{2}+\langle P\rangle^{2}).$ (25)
The results for using intensity as our network output during training and inference are shown in Fig. 14 for the XOR and handwritten digits classification problems using the same hyperparameters and circuit architecture considered in the main text. For the XOR problem the test accuracy remains at 100% while the test accuracy for the handwritten digits dataset decreases slightly to 97.08% which we note is still comparable to that achieved by the classical neural network trained for comparison in Section IV.2.2. The class probability heatmap for the XOR problem shows sharper definition of the decision boundary than the heatmap in Fig. 5(a) of the main text. However, a potential disadvantage of this approach is that regression for arbitrary real valued functions becomes less straightforward as the intensity measurement result is restricted to nonnegative numbers.
(a) Test set results for XOR classification problem.
(b) Test set results for handwritten digits classification problem.
Figure 14: Classification results with ONN using intensity as output logits for the XOR and handwritten digits classification problems.

### v3

@@ 4.2.1 Binary Classification
We generated data points for the last two problems using scikit-learn with the same Gaussian noise as in the XOR case. We generated 100 data points and used a 50/50 split of the data into training and test sets for all problems. The hyperparameters are given in the Supplemental Material [47]. We used a three-mode circuit with three layers for all binary classification problems and encode the inputs into the first two modes with the remaining mode being vacuum. The output logits were taken to be the expectation of the $Q$ quadrature of the first two modes of the circuit. The training data for each of the problems are shown in the Supplemental Material [47].
@@
Fig.4(d) shows the performance of the model when trained with photon loss where $\eta=0.5$ for the concentric circles problem. We can see that, as in the regression case, the model could still learn an appropriate nonlinear decision boundary in the presence of high losses. From the accompanying heatmaps provided in the Supplemental Material [47] we observed that, similar to the regression case larger displacements were applied in the lossy case. However, in contrast to the regression example all displacements throughout the network were higher in the lossy case, not just for the modes and layers closer to the output homodyne measurements.
@@
4.2.2 Multilabel Classification
In addition to the previous sections which demonstrate this ONN model can indeed solve non-trivial nonlinear learning problems we also investigate the performance of our network on more challenging and commonly used benchmarks of multilabel classification. We used three examples of multilabel classification: the Iris data set which is a commonly used benchmark [54, 55, 56, 57, 18, 16], the handwritten digits dataset considered in Refs. [58, 25] and the vowel recognition dataset which is widely used in benchmarking of physical neural networks [19, 59, 60, 61, 62, 17, 63]. The first two datasets are loaded via scikit-learn.
@@ 4.2.2 Multilabel Classification
The Iris dataset consists of 150 samples from three different species (classes) of Iris flower (setosa, virginica, and versicolor) with each sample being represented as a four dimensional vector with elements corresponding to petal width, petal length, sepal width and sepal length in cm [64]. We used 75 random samples for training and the rest for testing and normalized each input feature to be on the range $[0,1]$. We used a four mode circuit with three layers, leading to 42 trainable parameters, and the same hyperparameters used in the binary classification section. We encode the four features into the four modes as done in our previous experiments. The expectation value of the $Q$ quadrature of the first three modes are taken to be our models output logits. From the confusion matrix in the Supplemental Material [47] we can see that the test accuracy is $94.67\%$. The two species the model struggles slightly to differentiate are virginica and versicolor. This is in accordance with previous findings using this dataset in the literature [18] as the two species have significant overlap in feature space particularly with their sepal widths and sepal lengths as shown in the pair plot in the Supplemental Material [47].
@@
The handwritten digits dataset consists of 1797 gray scaled images of handwritten digits from 0 to 9 and was considered. For visualization purposes we show five example images in the Supplemental Material [47]. We used 60% of the data (1078 images) for training and 40% (719 images) for testing. The pixel values were normalized to be between $[0,1]$ and the images were flattened into a $64$ dimensional vector. For this problem we used a 64 mode circuit with four layers, leading to $8576$ trainable parameters, and took the expectation value of the $Q$ quadrature of the first 10 modes as the output logits of the model. We encode the 64 normalized pixel intensities into the 64 modes of the optical circuit. The hyperparameters used to train the ONN for this task are provided in the Supplemental Material [47]. We achieved a test accuracy of 98.19% as can be seen by examining the confusion matrix of the trained model provided in the Supplemental Material [47].
@@
We consider the vowel recognition dataset from Refs. [59, 61, 62, 63], which is a subset of the dataset used in Ref. [65], that consists of pronunciations of seven vowels by 37 female speakers where each vowel is characterized by 12 different formant frequencies. Specifically, each data point in the dataset is represented as a 12-dimensional vector containing the first three formants $F_{1},F_{2}$ and $F_{3}$ sampled at the steady state, $20\%$, $50\%$, and $80\%$ of the vowels pronunciation duration respectively. Each speaker pronounces each of the seven vowels leading to a dataset with 259 data points. We use a 12 mode circuit with three layers, leading to 270 trainable parameters, and the same hyperparameters as in the binary classification section with the learning rate set to $10^{-3}$. We use a random split of the data into 129 training samples and 130 testing samples. We normalize the 12 features such that they are all between $[0,1]$. From the confusion matrix in the Supplemental Material [47] we can see that the ONN achieves a test accuracy of 98.46%.
@@
5 Conclusion
We have proposed a method for constructing a linear optical neural network using laser interferometry and field displacements, which is the novelty claim of this paper. We demonstrated that this classical optical model displays remarkable resilience to photon loss which is a prominent source of imperfections in modern photonic integrated circuits [29]. The linear optical character of the circuit also makes it highly suitable for in situ training using various existing training protocols based on parameter shift rules [44], physical backpropagation [45], physics-aware training [61] or forward-only training [62, 66, 17, 16]. All this makes it particularly appealing for on-chip integrated optics implementations. We also investigated how our network performs on a variety of machine learning problems and its robustness to variance in the gradient estimates provided by parameter shift rules.
@@ 5 Conclusion
One straightforward quantum improvement would be to add an inline squeezer of the amplitude quadrature $Q$ before its measurement by homodyne detection during both training and inference. The corresponding quantum noise reduction will allow proportionally fewer measurements to determine the average $q=\langle Q\rangle$ consequently enabling faster inference and gradient estimation time. Note that just 10 dB of squeezing reduces the variance by ten times and that on-chip quantum homodyne detection has been demonstrated at a bandwidth of 9 GHz [67] and PIN photodiode detection bandwidth has been demonstrated to reach 265 GHz [68]. Another, albeit more complex, quantum improvement is swapping the single $M\times M$ interferometer with a set of $M$ single mode squeezers which are sandwiched between two $M\times M$ interferometers. This configuration, based on the singular value decomposition, is known to allow one to exactly mimic the functionality of fully connected linear layers within a classical neural network [22].
For future work one could explore how these networks can more closely approximate the expressivity of classical neural networks. For example, classical neural networks inherently have a feed forward structure in which outputs from one layer are composed with the nonlinear transformation (activation function) in the next layer leading to the network learning hierarchical and distributed representations of the inputs at each layer [38]. An interesting example of such feedforward (though a nonlinear optical one, using field encoding rather than parameter encoding) can be found in Ref. [17]. Furthermore, the existence of universal function approximation theorems for classical neural networks [69, 70] provides a strong theoretical basis for their expressivity and possible extensions of these theorems to our work merits further investigation. Lastly, extensions to implementing convolutional, recurrent and transformer layers are also important next steps.
@@
[…]
Funding OP was supported by U.S. National Science foundation grants OSI-2531569 [NQVL-QCAP NQVL:QSTD:Design: Quantum Computing Applications of Photonics (QCAP)], PHY-2514971, and ECCS-2530171. JMS acknowledges support from the National Science Foundation via DMR-2204312.
[…]
Acknowledgment We thank Benjamin Scellier, Maria Schuld and Catalina Albornoz for useful discussions.
[…]
Disclosures The authors declare no conflicts of interest.
[…]
Supplemental document See Supplemental 1 for supporting content.